# Coat stiffening explains the consensus pathway of clathrin-mediated endocytosis

### v1

@@ IV.1 Energy-based model
The second integral in Eq. (6) describes the polymerization and bending energies of the clathrin coat, where $\mu$ is the polymerization energy density, $\kappa_{\mathrm{c}}$ is the coat bending rigidity, and $H_{\mathrm{c}}$ is the preferred curvature of the coat. The coat energy is integrated only over the coated membrane parts. The preferred coat curvature is related to the preferred radius of the clathrin coat by $H_{\mathrm{c}}=1/R_{\mathrm{c}}$. The last term is the line tension energy, where $\zeta$ is the line tension. The line tension energy could be due to unsaturated clathrin bonds, additional protein binding at the edge of the clathrin coat or actin pushing on the edges.
The two integrals in principle should be evaluated over different neutral surfaces, namely the ones of membrane and coat, respectively, which have a typical distance of 15 nm from each other due to the gap layer of adaptor proteins and the finite thicknesses of the two layers. In order to formulate a transparent theory in the spirit of thin shell and surface Hamiltonian models, here we neglect this effect. This aspect of the model is closely related to the question of relevant model parameters; for example, the reported high values for coat rigidity typically apply for the composite including the gap, and not necessarily for the clathrin coat alone [27], as we will discuss below.
@@
where we used the membrane excess area $\Delta A=(1-\cos\theta)A/2$. The first term does not depend on size due to the conformal invariance of the bending Hamiltonian without spontaneous curvature. From Fig. 2F, we know the consensus pathway of invagination. Based on this, we can now predict the different energy contributions in absolute terms by using the coat area $A=2\pi R^{2}(1-\cos\theta)$ and the coat edge length $\mathcal{E}=2\pi R\sin\theta$ on Eq. (7).
@@
Coat radius $R_{\mathrm{c}}$ $32.5-50\,\mathrm{nm}$ [7, 26] $40\,\mathrm{nm}$ [17]
@@
We first note that both the membrane bending energy (Fig. 3A) and the tension energy (Fig. 3B) monotonically increase as a function of $\theta$. In contrast, the polymerization energy of the coat becomes more negative for increasing $\theta$ as the coat grows in size (Fig. 3C), in agreement with the general notion that release of free energy by polymerization is the main driving force for invagination. Because invagination starts with finite edge length, the line tension in Fig. 3D is approximately constant until the coat is halfway invaginated, before it rapidly decreases over the rest of the time course. This implies that line tension does not play a driving role during the initial phases of invagination. Later it could become important for invagination, as known for vesicle formation from phase-separated membrane domains [32, 33], multilayered membranes [34] and sometimes discussed in the context of CME [29, 25, 35], but our results show that the corresponding energies are rather small compared to the other contributions in Fig. 3. By far the largest energetic contribution comes from bending the clathrin coat (Fig. 3E). Because $\kappa_{c}$ has been determined in independent experiments, there is no doubt that this value will be very high. We see that a shallow minimum is caused by the coat approaching its preferred curvature, which decreases the energy; because at the same time area is still increasing, the energy then rises again. For large enough values of the preferred coat radius $R_{\mathrm{c}}$, the coat energy monotonically decreases (Fig. 3F). From these plots we conclude that the energetic competition between coat polymerization energy and coat rigidity strongly dominates the invagination pathway.
@@ IV.2 Total energy for static coat properties
Our goal is to predict the dominant physical mechanism for coat invagination and we assume that the overall invagination pathway has to be characterized by a well-defined downhill total energy landscape. Moreover, we expect that the flat state $(\theta=0)$ is characterized by a zero or negative total energy, so that flat growth is possible. Following our observation that polymerization energy and the stiffness of the coat are the dominant factors in coat invagination, we now investigate the effect of these two parameters. In Fig. 4A and B we plot the total energy as a function of $\theta$, given by Eq. (7), for different polymerization energies $\mu$ and coat rigidities $\kappa_{\mathrm{c}}$, respectively. We see that in all considered cases the total energy is positive over a significant part of the domain. Fig. 4A suggests that once started, a large enough polymerization energy could drive the invagination process over the whole lifetime, but that still the initial energy, associated with flat coat assembly, will be positive and hence make flat coat assembly at static parameter values unlikely. Fig. 4B indicates that the coat rigidity always renders the total energy positive for small values of $\theta$. We conclude that the large positive value of the coat bending energy for small values of $\theta$ makes flat clathrin coat assembly energetically costly and therefore potentially impossible. Thus the system in the initial stages should rather be characterized by small coat bending rigidity. On the other hand, in the late stages the coat bending rigidity should be high in order to enforce the preferred curvature. This suggests that coat bending rigidity might be time-dependent.
@@ IV.3 Total energy for dynamic coat properties
with $\alpha=\kappa_{\mathrm{c}}A\theta/k_{\mathrm{i}}$. In order to find an expression for $\partial\theta/\partial t$, we use the underlying dynamic assumption of the CoopCM, namely that the area $A$ of the clathrin coats grows along the edge $\mathcal{E}$ by addition of new triskelia with the growth speed $k_{\mathrm{on}}$ [16]
@@
with the rate of growth $k_{\mathrm{g}}=12\gamma k_{\mathrm{on}}/(8\gamma^{2}R_{0}^{2}-1)$, which carries the unit of $1\,\mathrm{s^{-1}}$. Now we can put everything together and obtain
@@
Figure 5: Total energy for time-varying model parameters. (A) The total energy including all energy contributions as predicted by the energetic model according to Eqs. (16-17) using the consensus invagination pathway of Fig. 2F. The used parameter values are indicated either in the figure legend or in Tab. 2. (B) The total energy including all energy contributions as predicted by the energetic model according to Eqs. (16-17) using the consensus invagination pathway of Fig. 2F, the pathway with the smallest pits (cf. Fig. 2A) and the pathway with the largest pits (cf. Fig. 2C). Parameter values used are indicated in Tab. 2. (C) Flat-to-curved transition predicted for $R_{\mathrm{c}}$, $R_{0}$, $k_{\mathrm{g}}$ and $k_{\mathrm{i}}$. Inset: The clathrin coat undergoes a flat-to-curved transition if $\xi\leq 1$.
The particular form of the second term of the prefactor suggests the assumption that we can neglect all energetic contributions except of the polymerization energy and coat rigidity on the right hand side of Eq. (14). When all energy contributions are normalized by $\kappa_{\mathrm{c}}$, all energy contributions except for the polymerization energy and the coat bending energy are rendered negligible, compare Fig. 3. Since the coat polymerization energy and the coat rigidity are the relevant quantities that drive coat invagination, we ask under which assumption we can get the CoopCM from these two terms. Therefore, we only consider the polymerization and coat energy in Eq. (7) and assume $\mu(H)$ and $\kappa(H)$ to depend on the degree of invagination. From Eq. (14) we obtain
@@
We note that we assume that the coat area $A$ grows by addition of triskelia over the edge of the coat, independent of H. In order to bring Eq. (15) to the same functional form as the CoopCM, given by Eq. (1), and to avoid any internal inconsistency within the model framework, the simplest possible assumption is that the polymerization energy increases linearly with the curvature. Moreover, we must assume that the clathrin coat stiffens with an exponent of $3$. We then have
@@
The assumption of an initially weak coat seems to be required to allow initial curvature generation. Later, the increase in this value reinforces the invagination because it forces the system to adapt the inherent curvature of the mature lattice.
@@
Using the fitted value of $R_{0}=63.2\,\mathrm{nm}$ and the polymerization energy $\mu=0.56\,\mathrm{mJ\,m^{-2}}$, we predict $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$ from Eq. (19). Compared to the typical value of coat rigidity in Tab. 2, our predicted value of $\kappa_{\mathrm{c}}$ is too small by a factor of 6 to 7, likely due to simplifications in our theory and the assumed functional form of Eqs. (16-17). In particular, a different curvature scale instead of $H_{\mathrm{c}}$ or higher order terms could contribute to coat stiffening. However, it is also possible that the experimentally determined value of $\kappa_{\mathrm{c}}$ is too high, especially because it usually includes the contributions of the mechanics of the gap layer [27]. In addition, using the fitted value of $\gamma=0.0110\,\mathrm{nm^{-1}}$, the coat rigidity of $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$ and the coat radius of $R_{\mathrm{c}}=40\,\mathrm{nm}$ we predict $k_{\mathrm{g}}/k_{\mathrm{i}}=5.5$ from Eq. (20). Our predicted value for $k_{\mathrm{g}}/k_{\mathrm{i}}$ indicates that the invagination of the clathrin coat is slower compared to growth, which makes intuitive sense and suggests that the coat closes due to growth rather than bending.
In Fig. 5A we plot Eq. (7) with varying polymerization energy and coat rigidity according to Eqs. (16-17) for the parameters in Tab. 2 and $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$. For both parameter combinations, the total energy is negative and monotonically decreasing as a function of $\theta$, as it should be in a physical description of CME. The result justifies our approach and suggests that polymerization energy and coat rigidity indeed dynamically increase during coat invagination.
@@
We finally predict the time course of the invagination of the clathrin coat. The flat-to-curved transition occurs when the area of the coat at the transition, $\pi R_{\mathrm{T}}^{2}$, is smaller than the final coat area, $4\pi R_{0}^{2}$, i.e., $\pi R_{\mathrm{T}}^{2}\leq 4\pi R_{0}^{2}$. Otherwise the clathrin coat grows flat. From the inequality, we deduce a dimensionless parameter $\xi=R_{\mathrm{T}}/(2R_{0})$ that predicts the flat-to-curved transition if $\xi\leq 1$, and flat growth otherwise (Fig. 5C, inset). We connect the invagination radius $R_{\mathrm{T}}={2k_{\mathrm{g}}\kappa_{\mathrm{c}}}/{(k_{\mathrm{i}}\mu R_{\mathrm{c}})}$, to the energetic parameters by using Eq. (20) and the final coat radius $R_{0}=\sqrt{{12\kappa_{\mathrm{c}}}/{\mu}}$ to the energetic parameters by using Eq. (19). Using these expressions on the condition of the flat-to-curved transition, we can thus relate the time course of invagination of the clathrin coat to the kinetic parameters, $k_{\mathrm{g}}/k_{\mathrm{i}}\leq 12R_{\mathrm{c}}/R_{0}$, which is illustrated in Fig. 5C. We conclude that the time course of invagination, parameterized by coat growth and invagination is determined by two length scales, the patch radius at which the flat-to-curved transition occurs and the final pit radius.
@@ V Discussion
In this work, we investigated the invagination pathway of clathrin coats. We combined our recently developed CoopCM with datasets from different cell lines and high-resolution imaging techniques (ET [13], HS-AFM [17] and SRM [16]). We found that the CoopCM gives excellent fits to the various datasets. Comparing the trajectories of the different datasets by means of the CoopCM, we found very similar invagination behavior. We concluded that clathrin coats follow a consensus invagination pathway described by the CoopCM. The model suggests that the dynamics of invagination is strongly determined by cooperative generation of curvature in the clathrin lattice. One very important result from both model and experiments is that a flat-to-curved transition occurs at finite coat area.
We then combined the consensus pathway and the CoopCM with an energetic model for the composite of membrane and coat. Our main result was that the energy related to coat stiffness dominates all other contributions. Moreover the flat-to-curved transition implies that the invagination process is not possible for energetic reasons if the coat stiffness is always at the high values reported experimentally for the later stages of the pits. Therefore, we conclude that coat stiffness has to start at low values and to dynamically increase during invagination. We showed that the CoopCM can be derived from the energetic model when assuming that the coat polymerization energy increases linearly with coat curvature and the clathrin coat stiffens with a power law exponent of three with coat curvature.
@@
We note that a power law exponent of three also relates bending stiffness and thickness of thin sheets, suggesting that effective thickness might grow linear with curvature. In practise, this viewpoint of continuum mechanics might however be too naive, because clathrin lattices have very specific discrete architectures. Rather, this prediction of our theory should be tested experimentally. Our prediction that the polymerization energy of the clathrin coat increases linearly with the curvature of the coat is consistent with experiments and theory predicting that clathrin assembly is stabilized on curved surfaces and that clathrin triskelia in curved coats contain more energy than in flat clathrin lattices [20, 37]. In fact, the possibility of a curvature-dependent polymerization energy has been discussed previously [30]. Possible driving factors for coat stiffening and increase of the effective thickness on a more microscopic level are accumulation of phosphorylated clathrin light chains [18], conformational changes in clathrin triskelia [22], rearrangements within the coat or mending of lattice defects [23], filling up lattice vacancies [24], release of elastic energy within the coat [17], increasing the density of the clathrin coat [38], solidification of the coat [29, 39], changing clathrin to AP2 adapter ratio during coat assembly [14], the organized binding of clathrin triskelia to adaptor clusters [40], mechanics of the coupling between membrane and clathrin coat through adapter proteins [27], and the stabilization of membrane curvature by clathrin in a ratchet-like manner [21].
@@
The stiffening of the clathrin coat could be complemented by an increasing line tension $\zeta$, for example by polymerizing actin pushing from the periphery of the coat [9, 10]. However, our data also shows that this might be dispensable under normal conditions, in line with previous experiments [43].
It will require future work to dissect how exactly the coat stiffening occurs. In principle this could happen in a continuous way, similar to releasing elastic energy within the coat [17], or in a discrete or ratchet-like way, for example due to conformational changes in the clathrin triskelia, rearrangements within the clathrin coat [22, 23], or by filling up lattice vacanices [24]. Therefore, experimentally the CoopCM could be tested by studying the thickness, the density or the mobility of single triskelia within the clathrin coat as a function of time.

### v2

@@ IV.1 Energy-based model
The second integral in Eq. (6) describes the polymerization and bending energies of the clathrin coat, where $\mu$ is the polymerization energy density, $\kappa_{\mathrm{c}}$ is the coat bending rigidity, and $H_{\mathrm{c}}$ is the preferred curvature of the coat. The coat energy is integrated only over the coated membrane parts. The preferred coat curvature is related to the preferred radius of the clathrin cage by $H_{\mathrm{c}}=1/R_{\mathrm{c}}$. The last term is the line tension energy, where $\zeta$ is the line tension. The line tension energy could be due to unsaturated clathrin bonds, additional protein binding at the edge of the clathrin coat or actin pushing on the edges.
In principle, the two integrals should be evaluated over different neutral surfaces, namely the ones of membrane and coat, respectively, which have a typical distance of 15 nm from each other due to the gap layer of adaptor proteins and the finite thicknesses of the two layers. In order to formulate a transparent theory in the spirit of thin shell and surface Hamiltonian models, here we neglect this effect. This aspect of the model is closely related to the question of relevant model parameters; for example, the reported high values for coat rigidity typically apply for the composite including the gap, and not necessarily for the clathrin coat alone [28], as we will discuss below.
@@
where we used the membrane excess area $\Delta A=(1-\cos\theta)A/2$. The first term does not depend on size due to the conformal invariance of the bending Hamiltonian without spontaneous curvature. From Fig. 2H, we know the consensus pathway of invagination. Based on this, we can now predict the different energy contributions in absolute terms by using the coat area $A=2\pi R^{2}(1-\cos\theta)$ and the coat edge length $\mathcal{E}=2\pi R\sin\theta$ on Eq. (7).
@@
Cage radius $R_{\mathrm{c}}$ $32.5-50\,\mathrm{nm}$ [7, 27] $40\,\mathrm{nm}$ [17]
@@
We first note that both the membrane bending energy (Fig. 3A) and the tension energy (Fig. 3B) monotonically increase as a function of $\theta$. In contrast, the polymerization energy of the coat becomes more negative for increasing $\theta$ as the coat grows in size (Fig. 3C), in agreement with the general notion that release of free energy by polymerization is the main driving force for invagination. Because invagination starts with finite edge length, the line tension in Fig. 3D is approximately constant until the coat is halfway invaginated, before it rapidly decreases over the rest of the time course. This implies that line tension does not play a driving role during the initial phases of invagination. Later it could become important for invagination, as known for vesicle formation from phase-separated membrane domains [33, 34], multilayered membranes [35] and sometimes discussed in the context of CME [30, 25, 36], but our results show that the corresponding energies are rather small compared to the other contributions in Fig. 3. By far the largest energetic contribution comes from bending the clathrin coat (Fig. 3E). Because $\kappa_{c}$ has been determined in independent experiments, there is no doubt that this value will be very high. We see that a shallow minimum is caused by the coat approaching its preferred curvature, which decreases the energy; because at the same time area is still increasing, the energy then rises again. For large enough values of the preferred cage radius $R_{\mathrm{c}}$, the coat energy monotonically decreases (Fig. 3F). From these plots we conclude that the energetic competition between coat polymerization energy and coat rigidity strongly dominates the invagination pathway.
@@ IV.2 Total energy for static coat properties
Our goal is to predict the dominant physical mechanism for coat invagination and we assume that the overall invagination pathway has to be characterized by a well-defined downhill total energy landscape. Moreover, we expect that the flat state $(\theta=0)$ is characterized by a zero or negative total energy, so that flat growth is possible. Following our observation that polymerization energy and the stiffness of the coat are the dominant factors in coat invagination, we now investigate the effect of these two parameters. In Figs. 4A and B we plot the total energy as a function of $\theta$, given by Eq. (7), for different polymerization energies $\mu$ and coat rigidities $\kappa_{\mathrm{c}}$, respectively. We see that in all considered cases the total energy is positive over a significant part of the domain. Fig. 4A suggests that once started, a large enough polymerization energy could drive the invagination process over the whole lifetime, but that still the initial energy, associated with flat coat assembly, will be positive and hence make flat coat assembly at static parameter values unlikely. Fig. 4B indicates that the coat rigidity always renders the total energy positive for small values of $\theta$. We conclude that the large positive value of the coat bending energy for small values of $\theta$ makes flat clathrin coat assembly energetically costly and therefore potentially impossible. Thus the system in the initial stages should rather be characterized by small coat bending rigidity. On the other hand, in the late stages the coat bending rigidity should be high in order to enforce the preferred curvature. This suggests that coat bending rigidity might be time-dependent.
@@ IV.3 Total energy for dynamic coat properties
with $\alpha=\kappa_{\mathrm{c}}A\theta/k_{\mathrm{i}}$. In order to find an expression for $\partial\theta/\partial t$, we use the underlying dynamic assumption of the CoopCM, namely that the area $A$ of the clathrin coat grows along the edge $\mathcal{E}$ by addition of new triskelia with the growth speed $k_{\mathrm{on}}$ [16]
@@
with the rate of growth $k_{\mathrm{g}}=12\gamma k_{\mathrm{on}}/(8\gamma^{2}R_{0}^{2}-1)$, which carries the unit of $1\,\mathrm{s^{-1}}$. Although the expansion in Eq. (13) is formally valid only for small values of $\theta$, the functional form is supported by an analysis of the invagination of clathrin coats as a function of pseudotime (the mapped frequency of invagination) [16]. The description deviates only for the late stages of coat assembly, where vesicle scission occurs. We therefore extrapolate the behavior as stated in Eq. (13) to the full domain of $\theta$. Now we can put everything together and obtain
@@
Figure 5: Total energy for time-varying model parameters. (A) The total energy including all energy contributions as predicted by the energetic model according to Eqs. (16-17) using the consensus invagination pathway of Fig. 2H. The used parameter values are indicated either in the figure legend or in Tab. 2. (B) The total energy including all energy contributions as predicted by the energetic model according to Eqs. (16-17) using the consensus invagination pathway of Fig. 2H, the pathway with the smallest pits (cf. Fig. 2A) and the pathway with the largest pits (cf. Fig. 2C). Parameter values used are indicated in Tab. 2. (C) Flat-to-curved transition predicted for $R_{\mathrm{c}}$, $R_{0}$, $k_{\mathrm{g}}$ and $k_{\mathrm{i}}$. Inset: The clathrin coat undergoes a flat-to-curved transition if $\xi\leq 1$.
The particular form of the second term of the prefactor ($1/(A\kappa_{\mathrm{c}})$) suggests the following assumptions on the right hand side of Eq. (14): When all energy contributions are normalized by $\kappa_{\mathrm{c}}$, all energy contributions except for the polymerization energy and the coat bending energy are rendered negligible (cf. Fig. 3). Since the coat polymerization energy and the coat rigidity are the relevant quantities that drive coat invagination, we ask under which assumption we can get the CoopCM from these two terms. Therefore, we only consider the polymerization and coat energy in Eq. (7) and assume $\mu(H)$ and $\kappa(H)$ to depend on the degree of invagination. From Eq. (14) we obtain
@@
Here we assume once again that the coat area $A$ grows by addition of triskelia over the edge of the coat, independent of H. In order to bring Eq. (15) to the same functional form as the CoopCM, given by Eq. (1), and to avoid any internal inconsistency within the model framework, the simplest possible assumption is that the polymerization energy increases linearly with the curvature. Moreover, we must assume that the clathrin coat stiffens with an exponent of $3$. We then have
@@
The assumption of an initially weak coat seems to be required to allow initial curvature generation. Later, the increase in this value reinforces the invagination because it forces the system to adapt to the inherent curvature of the mature lattice.
@@
Using the fitted value of $R_{0}=63.2\,\mathrm{nm}$ and the polymerization energy $\mu=0.56\,\mathrm{mJ\,m^{-2}}$, we predict $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$ from Eq. (19). Compared to the typical value of coat rigidity in Tab. 2, our predicted value of $\kappa_{\mathrm{c}}$ is too small by a factor of 6 to 7, likely due to simplifications in our theory and the assumed functional form of Eqs. (16-17). In particular, a different curvature scale instead of $H_{\mathrm{c}}$ or higher order terms could contribute to coat stiffening. However, it is also possible that the experimentally determined value of $\kappa_{\mathrm{c}}$ is too high, especially because it usually includes the contributions of the mechanics of the gap layer [28]. In addition, using the fitted value of $\gamma=0.0110\,\mathrm{nm^{-1}}$, the coat rigidity of $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$ and the cage radius of $R_{\mathrm{c}}=40\,\mathrm{nm}$ we predict $k_{\mathrm{g}}/k_{\mathrm{i}}=5.5$ from Eq. (20). We note that due to the functional form of Eq. (9), the dynamics of $\theta$, given by Eq. (12), will only scale the curvature evolution equation ($\partial H/\partial\theta$). Therefore, the connection between the energetic model and the CoopCM remains valid up to a scaling function even if a higher order expansion in $\theta$ in Eq. (13) or a different model for coat growth than Eq. (10) is used. In Fig. 5A we plot Eq. (7) with varying polymerization energy and coat rigidity according to Eqs. (16-17) for the parameters in Tab. 2 and $\kappa_{\mathrm{c}}=45\,\mathrm{k_{B}T}$. For both parameter combinations, the total energy is negative and monotonically decreasing as a function of $\theta$, as it should be in a physical description of CME. The result justifies our approach and suggests that polymerization energy and coat rigidity indeed dynamically increase during coat invagination.
@@
We finally predict the invagination pathway of the clathrin coat. The flat-to-curved transition occurs when the area of the coat at the transition, $\pi R_{\mathrm{T}}^{2}$, is smaller than the final coat area, $4\pi R_{0}^{2}$, i.e., $\pi R_{\mathrm{T}}^{2}\leq 4\pi R_{0}^{2}$. Otherwise the clathrin coat grows flat. From the inequality, we deduce a dimensionless parameter $\xi=R_{\mathrm{T}}/(2R_{0})$ that predicts the flat-to-curved transition if $\xi\leq 1$, and flat growth otherwise (Fig. 5C, inset). We connect the invagination radius $R_{\mathrm{T}}={2k_{\mathrm{g}}\kappa_{\mathrm{c}}}/{(k_{\mathrm{i}}\mu R_{\mathrm{c}})}$, to the energetic parameters by using Eq. (20) and the final coat radius $R_{0}=\sqrt{{12\kappa_{\mathrm{c}}}/{\mu}}$ to the energetic parameters by using Eq. (19). Using these expressions on the condition of the flat-to-curved transition, we can thus relate the invagination pathway of the clathrin coat to the kinetic parameters, $k_{\mathrm{g}}/k_{\mathrm{i}}\leq 12R_{\mathrm{c}}/R_{0}$, which is illustrated in Fig. 5C. We conclude that the invagination pathway, parameterized by coat growth and invagination is determined by two length scales, the patch radius at which the flat-to-curved transition occurs and the final pit radius.
@@ V Discussion
In this work, we investigated the invagination pathways of clathrin coats from different angles, combining the analysis of experimental data with modelling. The analyzed data sets were acquired with very different methods, namely ET [13], HS-AFM [17] and SRM [16], but all of them are high resolution spatial data that suggest a spherical cap shape of the clathrin pits and similar changes in curvature during invagination. The spherical cap assumption is further supported by a detailed analysis of the SRM data with a maximum-likelihood model [26]. Since the spherical cap matched the shape of the majority of clathrin coats in the SRM data [16], and only some sites exhibit asymmetric and irregular deformations, we believe that the spherical cap assumption is well suited to describe the coat shape in our theoretical model.
The geometrical time courses of the invaginations seem to be fitted well across imaging techniques and cell types by the recently introduced cooperative curvature model (CoopCM) [16], as shown in Fig. 2A-F. Comparing the trajectories of the different datasets by means of the CoopCM after correcting for experimental uncertainties, we found very similar invagination behavior (Fig. 2G). We concluded that clathrin coats follow a consensus invagination pathway described by the CoopCM (Fig. 2H). We note, however, that ultimately the CoopCM is derived phenomenologically and in principle other descriptions might exist, although, to date, we know of no other suitable candidate. In particular, a linear growth model was tested before to give worse results [16]. The CoopCM suggests that the dynamics of invagination is strongly determined by cooperative generation of curvature in the clathrin lattice. In the future, this general conclusion should be validated by a more microscopic model.
One important result from our model is that a flat-to-curved transition occurs at finite coat area, which agrees with experimental observations [13, 14, 16]. While our model describes the averaged pathway of coat invagination, we cannot rule out that different populations contribute to the analyzed datasets. Subpopulations might also follow different pathways and mechanisms for coat invagination. Nevertheless, we concluded that the CoopCM defines a consensus pathway which can be used for an energetic analysis.
We then combined the consensus pathway and the CoopCM with an energetic model for the composite of membrane and coat. Our main result was that the energy related to coat stiffness dominates all other contributions. Moreover, within the framework of our interface model and assuming no further energetic contributions other than those of the membrane and the clathrin coat, our analysis suggests that the coat stiffness changes during coat invagination. When using a constant value for stiffness and taking the high values measured for pits in their late stages, our model predicts that coat invagination is energetically unfavorable. Our model therefore predicts that coat stiffness starts at low values and dynamically increases during invagination, in agreement with the emerging notion that clathrin coats are plastic and cooperative. Although there is no experimental evidence to date, we point out that additional unreported molecular mechanisms may drive clathrin coat invagination.
@@
We showed that the CoopCM can be derived from the energetic model when assuming that the coat polymerization energy increases linearly with coat curvature and the clathrin coat stiffens with a power law exponent of three with coat curvature. We note that a power law exponent of three also relates bending stiffness and thickness of thin sheets [38], suggesting that effective thickness might grow linear with curvature. In practise, this viewpoint of continuum mechanics might however be too naive, because clathrin lattices have very specific discrete architectures. Rather, this prediction of our theory should be tested experimentally. Our prediction that the polymerization energy of the clathrin coat increases linearly with the curvature of the coat is consistent with experiments and theory predicting that clathrin assembly is stabilized on curved surfaces and that clathrin triskelia in curved coats contain more energy than in flat clathrin lattices [20, 39]. In fact, the possibility of a curvature-dependent polymerization energy has been discussed previously [31]. Possible driving factors for coat stiffening and increase of the effective thickness on a more microscopic level are accumulation of phosphorylated clathrin light chains [18], conformational changes in clathrin triskelia [22], rearrangements within the coat or mending of lattice defects [23], filling up lattice vacancies [24], release of elastic energy within the coat [17], increasing the density of the clathrin coat [40], solidification of the coat [30, 41], changing clathrin to AP2 adapter ratio during coat assembly [14], the organized binding of clathrin triskelia to adaptor clusters [42], mechanics of the coupling between membrane and clathrin coat through adapter proteins [28], and the stabilization of membrane curvature by clathrin in a ratchet-like manner [21].
@@
The stiffening of the clathrin coat could be complemented by an increasing line tension $\zeta$, for example by polymerizing actin pushing from the periphery of the coat [9, 10]. However, our data also shows that this might be dispensable under normal conditions, in line with previous experiments [45]. Moreover, the scenario would still require that the clathrin coat stiffens during invagination, because otherwise flat clathrin coat assembly cannot be explained. However, analyzing this scenario is beyond the scope of our model and future work will therefore be required. Our model is also not suited to analyze the invagination of clathrin coats that do not assume a spherical shape as for example in yeast [46].
Our interface model describes the mean invagination pathway of the clathrin coat and does not consider deviations from the spherical cap shape or clathrin lattices that are not regularly assembled. It will require future work and more microscopic models to dissect whether these aspects can impact the energetics of coat invagination and how coat stiffening occurs microscopically. In principle coat stiffening could happen in a continuous way, similar to releasing elastic energy within the coat [17], or in a discrete or ratchet-like way, for example due to conformational changes in the clathrin triskelia, rearrangements within the clathrin coat [22, 23], or by filling up lattice vacanices [24]. Therefore, experimentally the CoopCM could be tested by studying the thickness, the density or the mobility of single triskelia within the clathrin coat as a function of time.